# More Sample-Efficient Tuning of Particle Accelerators with Bayesian Optimization and Prior Mean Models

### v2

@@ Appendix B Neural Network Models
The LCLS injector model is a fully connected neural network implemented in PyTorch [25] with nine hidden layers, including up to nodes and Exponential Linear Unit (ELU) activation functions (). Furthermore, six dropout modules () are added after the 2nd to 7th layer to help prevent overfitting.
The model was trained initially on IMPACT-T simulation results and then calibrated to more closely match the measured data. The simulation data set was generated using Xopt. In this case, to solve the calibration problem, we froze the weights and biases of the original neural network and added linear nodes to each input. These nodes have free offset and scaling factors, i.e., a weight and a bias, with a linear activation function. These unknowns are then determined by re-training the modified network on measured data in the usual way (stochastic gradient descent and back-propagation through the neural network model). This enables simultaneous determination of calibration factors for multiple inputs and retains interpretability over a non-linear approach.
@@
The neural network model employed to predict transmission in the ATLAS accelerator, which takes five quadrupole currents as input, consists of two hidden layers, each with units, and employs the hyperbolic tangent as the activation function. The output layer utilizes a sigmoid function, constraining the transmission values to the range.
The model was implemented using PyTorch, with the MSE loss function employed for training. For optimization, the Adam optimizer was used with a learning rate of . Additionally, a learning rate scheduler was applied to adaptively adjust the learning rate during training based on the performance improvement.
The primary objective of implementing this model was to demonstrate that a simple approach for the prior mean could yield benefits in Bayesian optimization, as detailed in this paper. The model was trained using a dataset comprising experimental data points obtained from a prior experiment involving a 14N ion beam.
Subsequently, the trained model was deployed and evaluated on a different ion species, 16O, for another experiment. This process showcased the model’s capability to effectively transfer knowledge between different ion beams, highlighting its versatility and generalizability. Different models were trained to different levels of fidelity by conducting the training up to various epochs. The respective model fidelity is reflected in the correlation coefficient and MAE listed in Table 2.
Appendix C Practical Implementation of Neural Network Priors for Online Control
While the mathematics of incorporating neural networks into GP models is straightforward, the practical implementation in the context of online accelerator control can be challenging. For example, as discussed in Appendix B, the LCLS NN surrogate model was trained using simulation data and there is a significant mismatch between the names and units of the simulation parameters and those of the control parameters at the real machine. Translating the inputs and outputs of this model required a number of transformation layers, shown in Fig. 9, to map control parameters adjusted during optimization to NN input parameters. To address this, we implemented a wrapper class in the LUME-Model package [18] which stores the input and output ordering and the required transformations to make the model applicable to the real machine.

### v3

@@ A Gaussian process [3] (GP) is a distribution of functions denoted as
The LCLS injector model is a fully connected neural network implemented in PyTorch [18] with nine hidden layers, including up to nodes and Exponential Linear Unit (ELU) activation functions (). Furthermore, six dropout modules () are added after the 2nd to 7th layer to help prevent overfitting. The model was trained initially on IMPACT-T simulation results and then calibrated to more closely match the measured data. The simulation data set was generated using Xopt within the parameter ranges listed in Table 2. To study how the model’s accuracy impacts its use as a prior mean model, additional low-accuracy variations, listed in Table 1, were generated by varying the amount of training time. In order to solve the calibration problem, we froze the weights and biases of the original neural network and added linear nodes to each input. These nodes have free offset and scaling factors, i.e., a weight and a bias, with a linear activation function. These unknowns are then determined by re-training the modified network on measured data in the usual way (stochastic gradient descent and backpropagation through the neural network model). This enables simultaneous determination of calibration factors for multiple inputs and retains interpretability over a non-linear approach. The final model was then deployed as a prior mean model in Xopt within the parameter ranges listed in table 3.
| Epochs | Correlation | MAE () |
| Parameter | Min. | Max. | Unit |
| Laser Spot Size | |||
| Laser Pulse Length | |||
| Bunch Charge | |||
| Solenoid 121 | |||
| Quadrupole 121 | |||
| Quadrupole 122 | |||
| Linac 300 Amplitude | |||
| Linac 300 Phase | |||
| Linac 400 Amplitude | |||
| Linac 400 Phase | |||
| Quadrupole 361 | |||
| Quadrupole 371 | |||
| Quadrupole 425 | |||
| Quadrupole 441 | |||
| Quadrupole 511 | |||
| Quadrupole 525 |
| Name | Min. | Max. | Unit |
| Solenoid 121 | |||
| Quadrupole 121 | |||
| Quadrupole 122 | |||
| Quadrupole 361 | |||
| Quadrupole 371 | |||
| Quadrupole 425 | |||
| Quadrupole 441 | |||
| Quadrupole 511 | |||
| Quadrupole 525 |
@@
The neural network model employed to predict transmission in the ATLAS accelerator, which takes five quadrupole currents as input, consists of two hidden layers, each with units, and employs the hyperbolic tangent as the activation function. The output layer utilizes a sigmoid function, constraining the transmission values to the range. The model was implemented using PyTorch, with the MSE loss function employed for training. For optimization, the Adam optimizer was used with a learning rate of . Additionally, a learning rate scheduler was applied to adaptively adjust the learning rate during training based on the performance improvement. The primary objective of implementing this model was to demonstrate that a simple approach for the prior mean could yield benefits in Bayesian optimization, as detailed in this paper. The model was trained using a dataset comprising experimental data points obtained from a prior experiment involving a 14N ion beam. Subsequently, the trained model was deployed and evaluated on a different ion species, 16O, for another experiment. This process showcased the model’s capability to effectively transfer knowledge between different ion beams, highlighting its versatility and generalizability. Different models were trained to different levels of fidelity by conducting the training up to various epochs. The respective model fidelity is reflected in the correlation coefficient and MAE listed in Table 4.
| Epochs | Correlation | MAE () |
Practical Implementation of Neural Network Priors for Online Control
While the mathematics of incorporating neural networks into GP models is straightforward, the practical implementation in the context of online accelerator control can be challenging. For example, as discussed above, the LCLS NN surrogate model was trained using simulation data and there is a significant mismatch between the names and units of the simulation parameters and those of the control parameters at the real machine. Translating the inputs and outputs of this model required a number of transformation layers, shown in Fig. 12, to map control parameters adjusted during optimization to NN input parameters. To address this, we implemented a wrapper class in the LUME-Model package [12] which stores the input and output ordering and the required transformations to make the model applicable to the real machine.
@@
Data Availability
The data sets generated during the current study are available from the corresponding author on reasonable request.
Acknowledgments
This work was funded by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. It was also supported by the U.S. Department of Energy, Office of Nuclear Physics under Contract No. DE-AC02-06CH11357. The presented research used the ATLAS facility, which is a DOE Office of Science User Facility.
Author contributions statement
Conceptualization, T.B., J.L.M., B.M., D.R., R.R., and A.L.E.; Data curation, T.B., J.L.M., C.X., and K.R.L.B.; Formal analysis, T.B. and J.L.M.; Funding acquisition, D.R., A.L.E., and B.M.; Investigation, T.B. and J.L.M.; Methodology, R.R., T.B., J.L.M., and A.L.E.; Software, R.R., C.X., K.R.L.B., T.B., and J.L.M.; Experiment, T.B., J.L.M., R.R., and A.L.E.; Supervision, D.R., A.L.E., R.R., and B.M.; Validation, T.B. and J.L.M.; Visualization, T.B. and J.L.M.; Writing—original draft, T.B., J.L.M., R.R., B.M., and A.L.E.; Writing—review and editing, T.B., J.L.M., R.R., D.R., B.M., and A.L.E. All authors have read and agreed to the published version of the manuscript.
Additional information
Competing Interests
The author(s) declare no competing interests.